OpenStudy (gabylovesyou):

What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

first you need the slope of the line that goes through those two points

jimthompson5910 (jim_thompson5910):

so you use this formula m = (y2 - y1)/(x2 - x1)

OpenStudy (gabylovesyou):

m = -6 , 6/5

jimthompson5910 (jim_thompson5910):

m = (y2 - y1)/(x2 - x1) m = (-3 - 3)/(1 - (-5)) ... ... ... m = ??

OpenStudy (gabylovesyou):

-1 lol

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now onto finding the y-intercept y = mx + b y = -1x + b ... plug in the slope you just found 3 = -1(-5) + b ... plug in either point (I'm picking the first point) 3 = 5 + b now solve for b

OpenStudy (gabylovesyou):

b = -2

jimthompson5910 (jim_thompson5910):

so the equation of the line that goes through the two points is y = -1x - 2 or just y = -x - 2

jimthompson5910 (jim_thompson5910):

now plug in y = 0 and solve for x

OpenStudy (gabylovesyou):

x = -2

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (gabylovesyou):

the answer to the question is -2 ?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (gabylovesyou):

ok thnk you ! :)

jimthompson5910 (jim_thompson5910):

if you were to graph y = -x - 2, you'll see it crosses the x axis at (-2,0)

jimthompson5910 (jim_thompson5910):

yw